If you invest $5,550.00 in an account for 5 years that earns an annual rate of 6.50% compounded quarterly, what is the future value?
Answer by rfer(16322)

FV=5550(1+0.065/4)^4*5
FV=5550(1.01625)^20
FV=5550*1.38
FV=$7659.00